Find the indicated terms of the geometric sequence with the given description.
The common ratio is 0.75 and the fourth term is 864. Find the first three terms.
first term

second term

third term

Found 2 solutions by Boreal, greenestamps:
Answer by Boreal(15235)

xn=ar^(n-1)
x4=864
864=a^(3/4)^3
864/(27/64)=2048
first term is 2048
second term is 1536
third term is 1152

Answer by greenestamps(13200)



The common ratio moving forward in the sequence is 0.75 = 3/4.

Working backward in the sequence, the common ratio will be 4/3.

4th term = 864
3rd term: 864*(4/3) = 4*284 = 1152
2nd term: 1152*(4/3) = 4*384 = 1536
1st term: 1536*(4/3) = 4*512 = 2048